OpenStudy (anonymous):

how do you solve 1/5 more than a number m is 2 1/4

OpenStudy (anonymous):

m+1/5 = 2 1/4

OpenStudy (anonymous):

it is like saying 1 more than 2 is 3

OpenStudy (anonymous):

you could write that in mathematical form as 1+2 =3

OpenStudy (anonymous):

I trust you can take it from there?

OpenStudy (anonymous):

2 1/4 =\[2\frac{1}{4}\] ?